# A Generalized Back-Door Criterion for Linear Regression

### v2

@@ 1 Introduction
To analyze causal structures, we need to make certain assumptions about them. For example, Wright1921 introduced path analysis and illustrated the structures using path diagrams. This method has evolved through several studies, including Haavelmo1943, Joreskog1970, and Goldberger1972, and is also known as linear structural equation modeling (SEM). It is now widely used to analyze linear causal relationships.
@@
Meanwhile, some previous studies have explored the identification of partial causal relationships. Pearl1998 proposed the single-door criterion for identifying direct effects, and Avin2005 and Shpitser2008 studied the identifiability of path-specific effects. Tian2012, Foygel2012, and Barber2022 discussed the identifiability of path coefficients by considering the paths that contribute to covariance. In addition to the studies mentioned above, various other previous works exist; however, conditions guaranteeing that a partial regression coefficient coincides with a causal effect not mediated by covariates have been identified only in cases where the direct or total effect is of interest. Consequently, the logical validity of the causal interpretation of partial regression coefficients, as mentioned at the beginning, remains insufficiently established. Given that regression analysis is widely used not only in academic research but also in various practical contexts, it is crucial to provide a theoretically sound basis for its interpretation.
@@
Unlike the direction of generalization taken in previous studies, the selective-door criterion focuses solely on preserving directed paths that remain unblocked by covariates. As a result, it provides a clear framework for identifying partial causal relationships.
@@
The remainder of this paper is organized as follows. Section 2 introduces key terms and propositions as preliminaries. Section 3 formulates the problem addressed in this study and presents the main theorem. It also discusses the mechanism of post-treatment bias. Section 4 provides illustrative examples. All proofs are presented in Section 5.
@@ 2.1 Model Setup
where $\bm{PA}(X_{i})\subset\bm{V}\setminus\{X_{i}\}$ denotes the set of the immediate causes (parents) of $X_{i}$. Here, $a_{ij}$ is a nonzero constant referred to as the path coefficient from $X_{j}$ to $X_{i}$, and $u_{1},\dots,u_{N}$ are error terms, interpreted as the immediate causes that cannot be represented by the elements of $\bm{V}$. Throughout this paper, the error terms are assumed to be mutually uncorrelated.
@@
Assumption (ii) ensures that each causal diagram associated with equation (2.1) is a DAG. If we adhere to the view that the causal relationship between two variables at a given point in time is unidirectional, it is natural for the data-generating process to take the form of a DAG. However, this assumption also implies that a causal ordering must be specified. In practice, it is often necessary to consider multiple plausible causal orderings and compare the analytical results obtained under each scenario.
@@ 2.2 Regression on the Population
$Y=\bm{X}^{\top}\bm{\beta}+e,$
where $\bm{\beta}$ is a vector and $e$ is a random variable satisfying $\mathrm{E}[\bm{X}e]=\bm{0}$. Then, it follows that $\bm{\beta}=\,\mathrm{E}[\bm{X}\bm{X}^{\top}]^{-1}\,\mathrm{E}[\bm{X}Y]$.
@@ 2.3 Causal Diagram and the Selective-Door Criterion
Although a graph corresponding to a linear structural equation model is often referred to as a path diagram, this paper uses the term causal diagram to emphasize that it represents a data-generating process. A general discussion of causal diagrams can be found in Pearl2009.
@@
Pearl1993 focused on back-door paths, which can bias causal relationships, and established the back-door criterion. This criterion for covariate selection provides a sufficient condition for identifying the total causal effect.
@@
Moreover, this paper introduces the selective-door criterion, which is obtained by relaxing the back-door criterion.
@@ 2.4 Controlled Total Effects
See Wright1921 and Alwin1975 for details on the causal effects defined above. In addition, this paper introduces a measure representing partial causal relationships.
@@
Then, we summarize the validity of the total effect and the controlled total effect as indicators of causal relationships. It follows from ancestral expansion with $\bm{S}=\{X_{j}\}$ that
$X_{i}=\tau_{ij}X_{j}+\sum_{X_{k}\in\bm{AN}(X_{i})\setminus\{X_{j}\}}\tau_{ik|do(\{X_{j}\})}u_{k}+u_{i}.$
Since the error terms are interpreted as unobserved causes that cannot be represented by the elements of $\bm{V}$, the variable $X_{j}$ cannot affect $X_{i}$ through them. Therefore, $\tau_{ij}$ can be interpreted as the overall causal effect of $X_{j}$ on $X_{i}$. More generally, we have
$X_{i}=\sum_{X_{j}\in\bm{S}}\tau_{ij|do(\bm{S}_{-j})}X_{j}+\sum_{X_{k}\in\bm{AN}(X_{i})\setminus\bm{S}}\tau_{ik|do(\bm{S})}u_{k}+u_{i},$
which implies that $\tau_{ij|do(\bm{S}_{-j})}$ represents the effect of $X_{j}$ on $X_{i}$ that does not operate through any element of $\bm{S}_{-j}$. Using the do-operator (Pearl2009), this quantity can be expressed as
$\tau_{ij|do(\bm{S}_{-j})}=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}|do(X_{j}=x_{j}),do(\bm{S}_{-j}=\bm{s}_{-j})].$
@@ 3.2 Main Results
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E})$ be the causal diagram associated with equation (2.1). For any distinct nodes $X_{i},X_{j}\in\bm{V}$ and any subset $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ satisfying the selective-door criterion relative to $(X_{j},X_{i})$, the coefficient $\beta_{ij}$ in equation (3.1) coincides with the controlled total effect $\tau_{ij|do(\bm{Z})}$.
@@
As corollaries, we obtain the following propositions, originally presented by Pearl1998.
@@
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E})$ be the causal diagram associated with equation (2.1). For any distinct nodes $X_{i},X_{j}\in\bm{V}$, suppose that a subset $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}\cup\bm{DE}(X_{i}))$ blocks every path from $X_{j}$ to $X_{i}$ except for the directed edge $X_{j}\rightarrow X_{i}$. Then, the coefficient $\beta_{ij}$ in equation (3.1) coincides with the direct effect of $X_{j}$ on $X_{i}$.
@@
This paper also provides sufficient conditions under which all regression coefficients can be interpreted as causal effects.
@@ 3.3 Post-Treatment Bias
$\gamma_{ik}\coloneqq\beta_{ik}-\tau_{ik|do(\bm{S}\setminus\{X_{k}\})},\quad X_{k}\in\bm{S}.$
@@ 4 Examples
By Theorem 3.1 and Proposition 3.4, we can easily confirm this equation coincides with ancestral expansion. That is, it does not include the regression coefficients in equation (4.4), but instead corresponds to the controlled total effects composed solely of the path coefficients in the true data-generating process (4.3).
@@ 5 Proofs
Let $\bm{X}$$_{+}$ denote $(1,\bm{X}^{\top})^{\top}$. It is sufficient to prove the existence of $\mathrm{E}[\bm{X}_{+}\bm{X}_{+}^{\top}]^{-1}$. Since $\bm{I}_{N}-\bm{A}$ is a lower unitriangular matrix and nonsingular, equation (2.3) has another form
@@
Define two random variables, $e_{i}$ and $e_{i}$, as follows:
$e_{i}$ $=$ $\sum_{X_{l}\in\bm{AN}(X_{i})\setminus\bm{S}}\tau_{il|do(\bm{S})}u_{l}+u_{i},$
$e^{*}_{i}$ $=$ $e_{i}-\gamma_{i0}-\sum_{X_{m}\in\bm{Z}^{*}}\gamma_{im}X_{m},$
@@
  • For any $X_{p}\in\bm{Z}^{*}_{1}$, there exits a back-door path from $X_{p}$ to $X_{i}$ that is unblocked by $\bm{S}\setminus\{X_{p}\}$.
@@
Here, we follow the notation used in the proof of Theorem 3.1; hence, we have $\bm{S}_{1}=\bm{X}^{*}\cup\bm{Z}^{*}_{2}$ and $\bm{S}_{2}=\bm{Z}^{*}_{2}$. The following equation follows from the proof.
@@
$=$ $\tilde{\gamma}_{i0}+\sum_{X_{p}\in\bm{Z}^{*}_{1}}\tilde{\gamma}_{ip}\left(X_{p}-\sum_{X_{q}\in\bm{Z}^{*}_{2}}\tau_{pq|do(\bm{S}_{2}\setminus\{X_{q}\})}X_{q}\right)+\tilde{e}$.
@@
$\gamma_{iq}$ $=$ $-\sum_{X_{p}\in\bm{Z}^{*}_{1}}\gamma_{ip}\tau_{pq|do(\bm{S}_{2}\setminus\{X_{q}\})},\quad X_{q}\in\bm{Z}^{*}_{2}.$
Acknowledgements

### v3

@@ 1 Introduction
To analyze causal structures, we need to make certain assumptions about them. For example, Wright1921 introduced path analysis and illustrated the structures using path diagrams. This method has evolved through several studies, including Haavelmo1943, Joreskog1970, Goldberger1972, and Bollen1987, and is also known as linear structural equation modeling (SEM). It is now widely used to analyze linear causal relationships.
@@
Meanwhile, some previous studies have explored the identification of partial causal relationships. Pearl1998 proposed the single-door criterion for identifying direct effects, and Avin2005 and Shpitser2008 studied the identifiability of path-specific effects. BritoandPearl2002, Tian2009, Foygel2012, Chen2016, and Barber2022 discussed the identifiability of path coefficients by considering the paths that contribute to covariance. In addition to the studies mentioned above, various other previous works exist; however, conditions guaranteeing that a partial regression coefficient coincides with a causal effect not mediated by covariates have been identified only in cases where the direct or total effect is of interest. Consequently, the logical validity of the causal interpretation of partial regression coefficients, as mentioned at the beginning, remains insufficiently established. Given that regression analysis is widely used not only in academic research but also in various practical contexts, it is crucial to provide a theoretically sound basis for its interpretation.
@@
Unlike the direction of generalization taken in previous studies, the selective-door criterion focuses solely on preserving directed paths that remain unblocked by covariates. As a result, it provides a clear framework for identifying partial causal relationships. Moreover, whereas Pearl1998 discussed regression coefficients that minimize the mean squared error under conditional distributions of linear Gaussian models, in this paper, without assuming any specific distribution for the error terms and by employing uncorrelatedness, I present general results for standard partial regression coefficients. This means that linear causal effects can be estimated by the least squares method without assuming independence or normality of the error terms in the data-generating process.
@@
The remainder of this paper is organized as follows. Section 2 introduces key terms and propositions as preliminaries. Section 3 formulates the problem addressed in this study and presents the main theorem. The section also discusses the mechanisms of post-treatment bias. Section 4 provides illustrative examples. All proofs are presented in Section 5.
@@ 2.1 Model Setup
where $\bm{PA}(X_{i})\subset\bm{V}\setminus\{X_{i}\}$ denotes the set of immediate causes (parents) of $X_{i}$. Here, $a_{ij}$ is a nonzero constant referred to as the path coefficient from $X_{j}$ to $X_{i}$, and $u_{1},\dots,u_{N}$ are error terms interpreted as aggregations of all causes not included in $\bm{V}$. In studies of linear SEM, error terms are often assumed to be mutually independent or normally distributed. This paper relaxes this assumption, assuming instead that they are uncorrelated.
@@
Assumption 2.1 ensures that each causal diagram associated with equation (2.1) is a DAG. If we adhere to the view that the causal relationship between two variables at a given point in time is unidirectional, it is natural for the data-generating process to take the form of a DAG. However, this assumption also implies that a causal ordering must be specified. In practice, it is often necessary to consider multiple plausible causal orderings and compare the analytical results obtained under each scenario.
@@ 2.2 Regression on the Population
$Y=\bm{X}^{\top}\bm{b}+e,$
where $\bm{b}$ is a vector and $e$ is a random variable satisfying $\mathrm{E}[\bm{X}e]=\bm{0}$. Then, it follows that $\bm{b}=\,\mathrm{E}[\bm{X}\bm{X}^{\top}]^{-1}\,\mathrm{E}[\bm{X}Y]$.
@@ 2.3 Causal Diagram and the Selective-Door Criterion
Unless specified otherwise, a directed edge is regarded as a directed path of length one. In the following, we introduce several terms.
@@
According to the standard definition (see Pearl2009), bidirected arrows are drawn between nodes whose error terms are dependent. However, in this paper, bidirected arrows are omitted when the corresponding error terms are uncorrelated. Next, we introduce the following concepts to visually identify the conditional relationships between variables.
@@
Pearl1993 focused on back-door paths, which can bias causal relationships, and established the back-door criterion. This criterion for covariate selection provides a sufficient condition for identifying total causal effects.
@@
Although the back-door criterion serves as a powerful framework for covariate selection, it suffers from a limitation in capturing partial causal effects originating from specific paths. To address this issue, we relax the first condition of the back-door criterion and propose a generalized version.
@@ 2.4 Controlled Total Effects
See Alwin1975 and Bollen1987 for details on the causal effects defined above. In addition, this paper introduces a measure representing partial causal relationships.
@@
Using the do-operator Pearl2009, this quantity can be interpreted as
$\tau_{ij|do(\bm{S})}=\frac{\partial}{\partial x_{j}}\,\mathrm{E}[X_{i}|do(X_{j}=x_{j}),do(\bm{S}=\bm{s})].$
@@
Note that, on the right-hand side of this equation, the coefficients $\tau_{ij|do(\bm{S}_{-j})}$ corresponding to any random variables in $\bm{S}\setminus\bm{AN}(X_{i})$ are equal to 0. Since equation (2.4) is derived solely through successive substitution operations in the linear structural equations, it can be regarded as an alternative representation of the data-generating process.
@@ 3.2 Main Results
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E})$ be the causal diagram associated with equation (2.1). For any distinct nodes $X_{i},X_{j}\in\bm{V}$ and any subset $\bm{Z}\subset\bm{V}\setminus\{X_{i},X_{j}\}$ satisfying the selective-door criterion relative to $(X_{j},X_{i})$, the coefficient $\beta_{ij}$ in equation (3.1) coincides with the $\bm{Z}$-controlled total effect $\tau_{ij|do(\bm{Z})}$.
@@
As corollaries, we obtain the following propositions.
Corollary 3.2.
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E})$ be the causal diagram associated with equation (2.1). For any distinct nodes $X_{i},X_{j}\in\bm{V}$, suppose that a subset $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}$ blocks every path between $X_{j}$ and $X_{i}$. Then, the coefficient $\beta_{ij}$ in equation (3.1) is equal to 0.
@@
Assume that Assumption 2.1 holds. Let $\bm{G}=(\bm{V},\bm{E})$ be the causal diagram associated with equation (2.1). For any distinct nodes $X_{i},X_{j}\in\bm{V}$, suppose that a subset $\bm{Z}\subset\bm{V}\setminus(\{X_{i},X_{j}\}\cup\bm{DE}(X_{i}))$ blocks every path between $X_{j}$ and $X_{i}$ except for the directed edge $X_{j}\rightarrow X_{i}$. Then, the coefficient $\beta_{ij}$ in equation (3.1) coincides with the direct effect of $X_{j}$ on $X_{i}$.
@@
Pearl1998 has shown that these results hold under certain assumptions for conditional regression coefficients, i.e., the linear best approximation coefficients under the conditional distribution. A contribution of the present study is to demonstrate that causal effects can be consistently estimated using ordinary least squares, even in the absence of assumptions regarding the independence or normality of the error terms in the data-generating process.
In many empirical analyses, it is not uncommon for all explanatory variables to be regarded as causes of interest. Thus, this paper also provides sufficient conditions under which all regression coefficients can be interpreted as causal effects.
@@ 3.3 Post-Treatment Bias
$\gamma_{ik}\coloneq\beta_{ik}-\tau_{ik|do(\bm{S}\setminus\{X_{k}\})},\quad X_{k}\in\bm{S}.$
@@ 4 Examples
By Theorem 3.1 and Proposition 3.5, we can easily confirm that it does not include the regression coefficients in equation (4.4), but instead corresponds to the controlled total effects composed solely of the path coefficients in the true data-generating process (4.3). This can also be confirmed by Proposition 2.14 and Proposition 2.5.
@@ 5 Proofs
Let $\bm{X}$$_{+}$ denote $(1,\bm{X}^{\top})^{\top}$. It is sufficient to prove the existence of $\mathrm{E}[\bm{X}_{+}\bm{X}_{+}^{\top}]^{-1}$. Since $\bm{I}_{N}-\bm{A}$, where $\bm{I}_{N}$ is the $N\times N$ identity matrix, is a lower unitriangular matrix and nonsingular, equation (2.3) has another form
@@
Define two random variables, $e_{i}$ and $e^{*}_{i}$, as follows:
$e_{i}$ $\coloneq\sum_{X_{l}\in\bm{AN}(X_{i})\setminus\bm{S}}\tau_{il|do(\bm{S})}u_{l}+u_{i},$
$e^{*}_{i}$ $\coloneq e_{i}-\gamma_{i0}-\sum_{X_{m}\in\bm{Z}^{*}}\gamma_{im}X_{m},$
@@
  • For any $X_{p}\in\bm{Z}^{*}_{1}$, there exists a back-door path from $X_{p}$ to $X_{i}$ that is unblocked by $\bm{S}\setminus\{X_{p}\}$.
@@
Here, we follow the notation used in the proof of Theorem 3.1; hence, we have $\bm{S}_{1}=\bm{Z}^{*}_{1}$ and $\bm{S}_{2}=\bm{X}^{*}\cup\bm{Z}^{*}_{2}$. The following equation follows from the proof.
@@
$=\tilde{\gamma}_{i0}+\sum_{X_{p}\in\bm{Z}^{*}_{1}}\tilde{\gamma}_{ip}\left(X_{p}-\sum_{X_{q}\in\bm{Z}^{*}_{2}}\tau_{pq|do((\bm{X}^{*}\cup\bm{Z}^{*}_{2})\setminus\{X_{q}\})}X_{q}\right)+\tilde{e}$.
@@
$\gamma_{iq}$ $=-\sum_{X_{p}\in\bm{S}_{1}}\gamma_{ip}\tau_{pq|do(\bm{S}_{2}\setminus\{X_{q}\})},\quad X_{q}\in\bm{Z}^{*}_{2}.$
Acknowledgments